OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y =____x +____

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

y=mx+b Soooo... 2=3(6)+b therefore 2=18+b -18 from both sides to get -16=b

OpenStudy (anonymous):

i don't have to solve it

OpenStudy (anonymous):

Oh... Oh well... :)

Directrix (directrix):

Use the point slope calculator from the other day. I'll see if I can find the URL

OpenStudy (anonymous):

ok btw i don't need to solve it

Directrix (directrix):

This is better: https://www.easycalculation.com/analytical/point-slope.php